### v1

# Crowd Comparative Reasoning: Unlocking Comprehensive Evaluations for LLM-as-a-Judge

Abstract
LLM-as-a-Judge, which generates chain-of-thought (CoT) judgments, has become a widely adopted auto-evaluation method. However, its reliability is compromised by the CoT reasoning’s inability to capture comprehensive and deeper details, often leading to incomplete outcomes. Existing methods mainly rely on majority voting or criteria expansion, which is insufficient to address the limitation in CoT. We propose Crowd-based Comparative Evaluation, which introduces additional crowd responses to compare with the candidate responses, thereby exposing deeper and more comprehensive details within the candidate responses. This process effectively guides LLM-as-a-Judge to provide a more detailed CoT judgment. Extensive experiments demonstrate that our approach enhances evaluation reliability, achieving an average accuracy gain of $6.7\%$ across five benchmarks. Moreover, our method produces higher-quality CoTs that facilitate judge distillation and exhibit superior performance in rejection sampling for supervised fine-tuning (SFT), referred to as crowd rejection sampling, thereby enabling more efficient SFT. Our analysis confirms that CoTs generated by ours are more comprehensive and of higher quality, and evaluation accuracy improves as inference scales.
Crowd Comparative Reasoning: Unlocking Comprehensive Evaluations for LLM-as-a-Judge
Qiyuan Zhang1thanks: Work partially done during the internship at Huawei Noah’s Ark Lab., Yufei Wang2, Yuxin Jiang3, Liangyou Li2, Chuhan Wu2, Yasheng Wang2, Xin Jiang2, Lifeng Shang2, Ruiming Tang2, Fuyuan Lyu4, Chen Ma1thanks: Corresponding Author. 1City University of Hong Kong, 2Huawei Noah’s Ark Lab, 3The Hong Kong University of Science and Technology (Guangzhou), 4McGill University & MILA qzhang732-c@my.cityu.edu.hk, wang.yufei1@huawei.com,
1 Introduction
With the prohibitive cost and limited scalability of human evaluation, LLM-as-a-Judge has emerged as a scalable framework for auto-evaluation (Chang et al., 2024; Li et al., 2024a, 2025). Given a task instruction and corresponding candidate responses, LLM-as-a-Judge (Zheng et al., 2023; Wang et al., 2024b; Wagner et al., 2024) employs CoT judgment to analyze granular quality details of the responses, ultimately deriving a final outcome. Despite advancements in techniques such as CoT reasoning (Saha et al., 2025; Zheng et al., 2023), specialized rubrics (Liu et al., 2023), and preference-aligned training datasets (Li et al., 2024b; Wang et al., 2024c), human evaluation remains the gold standard due to persistent limitations (Zeng et al., 2024a) in LLM-as-a-Judge. These limitations include biases (Park et al., 2024) in judgment and susceptibility to misleading context (Dubois et al., 2024a; Chen et al., 2024), which undermine the reliability of automated evaluation. One important yet overlooked reason is the quality of CoT reasoning hinges on the model’s ability to comprehensively compare nuanced details across responses. Our observation reveals high-quality judgments incorporate a thorough comparison of these details, while flawed reasoning tends to focus on limited details, leading to premature and incomplete outcomes. Therefore, enhancing the richness and comprehensiveness of CoT reasoning is essential to improve LLM-as-a-Judge.
Figure 1: An overview of our method. By evaluating the candidate responses A/B alongside the crowd responses, the resulting crowd judgment can be used as context to enrich the evaluation of A/B responses, leading to a more comprehensive CoT judgment.
Two commonly adopted strategies aim to address this issue: majority voting (Zhang et al., 2024; Mahan et al., 2024; DeepSeek-AI, 2024) and criteria expansion (Kim et al., 2024a; Liu et al., 2024; Hu et al., 2024a). The majority voting generates multiple judgments independently in parallel and aggregates these results through voting. It essentially leverages the randomness from temperature sampling to encourage detailed reasoning. However, this approach is passive and computationally expensive. In contrast, criteria expansion augments prompts with additional evaluation aspects, proactively guiding the model to consider more dimensions of quality. Yet, this strategy is response-unaware, failing to adapt the evaluation process to the unique details of each response. For instance, even if a response is rich with nuanced insights, incorporating a criterion like “accuracy” does little to prompt the LLM to identify the unique details of its reasoning. Consequently, neither approach effectively guides LLM-as-a-Judge to consistently produce nuanced, comprehensive CoT evaluations. This leads to a critical research question: how can we guide LLMs to engage in deeper, more detail-rich CoT reasoning during judgment?
In this work, we propose a novel crowd-based comparative evaluation (CCE) to address this challenge by enabling LLM-as-a-Judge to uncover valuable details, as depicted in Figure 1. Our approach is inspired by human evaluative behavior: humans merely compare candidates in isolation by also contrasting them against a broader crowd, thereby uncovering additional nuanced insights about each candidate. Building on this principle, CCE first gathers a set of alternative responses to the task instruction, referred to as crowd responses, and then compares each candidate response against these crowd responses to derive multiple crowd judgments. Throughout this process, the diversity of crowd responses serves as multiple evaluation anchors, revealing different layers of detail within the candidate responses. Based on this, CCE prompts the LLM-as-a-Judge to perform a more comprehensive and deeper overall CoT judgment.
CCE achieves a remarkable average improvement of $6.7\%$ across five evaluation benchmarks, including RewardBench, HelpSteer2, MTBench Human, JudgeBench and EvalBias. When applied to judge distillation, we find high-quality long CoT judgments produced by CCE have a higher efficiency for training a smaller judge model, especially enhancing bias robustness. Moreover, we extend CCE naturally to SFT rejection sampling, refer to crowd rejection sampling, where our approach serves as a quality signal to identify training-efficient samples from the response pool. This extension highlights the reliability and practicality of CCE. Finally, analysis experiments confirm that CCE promotes more comprehensive and deeper CoT reasoning and scales inference effectively.
2 Related Work
Human evaluation is typically regarded as the gold standard for evaluating LLM responses to intricate and open-ended instructions Chiang and Lee (2023); Elangovan et al. (2024). Nevertheless, due to its inherent limitations—being time-consuming, costly, and prone to variability Karpinska et al. (2021)—automated evaluation methods leveraging LLMs have gained prominence as scalable and cost-efficient alternatives. Unlike reward models that provide only scalar scores Wang et al. (2024a, b), LLM-as-a-Judge frameworks offer enhanced robustness and interpretability by producing detailed CoT rationales Li et al. (2024c); Gao et al. (2024).
Enhancing the performance of LLM-as-a-Judge has attracted significant attention, with many techniques proposed recently. One prominent approach involves fine-tuning pre-trained LLMs on task-specific datasets to better adapt them for judgment tasks Vu et al. (2024); Li et al. (2024b); Wang et al. (2024c); Kim et al. (2024b). Another line of research focuses on step-by-step methodologies, such as G-EVAL Liu et al. (2023), ICE-Score Zhuo (2024), and EvalPlanner Saha et al. (2025), which decompose complex evaluation tasks into granular components, thereby harnessing the reasoning capabilities of LLMs to streamline the evaluation process. Additionally, recent advances explore using LLMs to generate reasoning traces by designing domain-specific prompts and meticulously crafting components of CoT reasoning. These include constructing fine-grained scoring rubrics Zheng et al. (2023); Zeng et al. (2024b); Trivedi et al. (2024) and generating reference answers Zhang et al. (2025). Despite these efforts, the richness and comprehensiveness of CoT reasoning remain underexplored, leaving room for further advancements in improving LLM-as-a-Judge. While simple heuristics such as majority voting Badshah and Sajjad (2024); Verga et al. (2024) can mitigate this issue by improving the reliability and accuracy of evaluations, they often fall short in terms of efficacy and efficiency.
Figure 2: Pipeline of our proposed crowd-based comparative evaluation. For a given instance $(x,y^{A},y^{B})$, we first use the LLM to generate crowd responses $\left\{y^{i}|i\in\{C,D,E,...\}\right\}$ based on $x$. These responses are then compared with $y^{A}$ and $y^{B}$ to produce initial crowd judgments $\mathcal{J}$, which are subsequently refined into $\hat{\mathcal{J}}$ after selection and processing. Finally, $\hat{\mathcal{J}}$ are used as contextual input to evaluate the instance $(x,y^{A},y^{B})$.
3 Methodology
As illustrated in Figure 2, we propose a crowd-based comparative evaluation that elicits and integrates multiple crowd judgments before producing a final outcome. It consists of three core components: (1) Crowd Response and Judgment Generation, (2) Crowd Judgment Selection and Processing, and (3) Context-augmented Inference, which we will discuss in the following subsections. Furthermore, we distill the CoT judgments generated by CCE to train a judge and expand its application to an enhanced rejection sampling technique for SFT.
3.1 Problem Formulation
Supposing $\{y^{A},y^{B}\}$ denote two candidate responses generated by two assistants for a given task instruction $x$, Vanilla LLM-as-a-Judge $\mathcal{F}$ is prompted to provide a CoT-based judgment $j$ of $y^{A}$ and $y^{B}$, based on a specific set of evaluation criteria $s$ (e.g., correctness, coherence).
$j=\mathcal{F}(y^{A},y^{B}|x,s).$ (1)
The objective is to ensure that the $\mathcal{F}$ preference aligns closely with human evaluation. In pairwise comparisons, this alignment is quantified by measuring the accuracy relative to human labels.
3.2 Crowd Response and Judgment Generation
Based on the task instruction $x$, we first prompt the LLM to generate a set of $n$ synthetic crowd responses $\left\{y^{i}|i\in\{C,D,E,...\}\right\}$. To enhance the diversity of these responses, we can leverage multiple LLMs ranging from smaller models (e.g., Qwen2.5-0.5B-Instruct) to larger ones (e.g., Mistral-Nemo-Instruct-2407), along with varying temperature settings. Theoretically, more diverse responses can cover a wider range of scenarios. When compared with $y^{A}$ and $y^{B}$, these crowd responses emphasize different details of $\{y^{A},y^{B}\}$, offering a more comprehensive perspective and facilitating deeper reasoning. As Figure 2 demonstrated, crowd judgment digs the importance of “he”, where Response A subtly shifts the actor “he” onto the object “task” itself, thereby violating the instruction’s requirement to rewrite while preserving the concise original meaning. Then, we use it as context to reinforce the following CoT. This advantage surpasses that of criteria expansion, which cannot anticipate such details through pre-prompting.
For each synthetic $y^{i}$, $\mathcal{F}$ independently produces two crowd judgments, $j^{A}_{i}$ and $j^{B}_{i}$, by individually judging $y^{i}$ with $y^{A}$ and $y^{B}$, separately:
$j^{A}_{i}=\mathcal{F}(y^{A},y^{i}|x,s),\quad j^{B}_{i}=\mathcal{F}(y^{B},y^{i}|x,s).$ (2)
Formally, we collect a set of $2n$ crowd judgments:
$\mathcal{J}=\left\{j^{A}_{i},j^{B}_{i}\mid i\in\{C,D,E,...\}\right\}.$ (3)
While each judgment may not fully capture all details of the candidate responses, they together provide a richer pool of evidence about how $y^{A}$ and $y^{B}$ differ in nuanced ways.
3.3 Crowd Judgment Selection and Processing
After obtaining $\mathcal{J}$, the key stage lies in selecting and processing these judgments effectively. Random Selection is neither stable nor optimal, so we need better strategies for using crowd judgments.
To this end, we propose a simple yet effective method called Criticizing Selection. Specifically, we choose judgments based on their outcomes: for $j_{i}^{A}$, we keep those where A loses, and for $j_{i}^{B}$, those where B loses. Notably, our observation reveals judgments with a critical outcome tend to provide detailed and informative reasoning for the criticized response. For instance, Judge might point out how the criticized response confuses key concepts by elaborating on specific errors in the definition and citing relevant theoretical principles. In contrast, judgments favoring the winning response tend to be brief, where the Judge might simply say, “this answer is correct” without further analysis. We also explore two alternative outcome-based strategies: Praising Selection (choosing only judgments where A/B wins) and Balanced Selection (maintaining an equal split between A/B wins and losses). However, as shown in our analysis (Table 4), both strategies perform worse than Criticizing Selection. Additionally, to mitigate bias from the outcome distribution from crowd judgments, we introduce Outcome Removal, where an LLM rewrites $j_{i}$ to remove explicit outcome segments, ensuring a more neutral evaluation. After the selection and processing, we obtain $\hat{{\mathcal{J}}}$.
Notably, $j_{i}$ includes CoT judgments not only of the ($y^{A}$, $y^{B}$) but also of $y^{i}$. Our pilot study shows that removing the CoT segments about $y^{i}$ does not improve performance; therefore, we retain them to keep our approach simple.
3.4 Context-augmented Inference
The final judgment is derived by evaluating responses $y^{A}$ and $y^{B}$ conditioned on the instruction $x$, the criteria $s$, and the post-processed crowd judgments $\hat{\mathcal{J}}$:
$j^{\star}=\mathcal{F}(y^{A},y^{B}\mid x,s,\hat{\mathcal{J}}),$ (4)
where the prompt template is provided in Appendix A. Notably, we distill $\{j^{\star}\}$ for training a smaller judge, whose performance surpasses the judge distilled from $\{j\}$, as demonstrated in Table 2. It proves that higher-quality CoT judgment has better distillation efficiency.
3.5 Extensive Application–Crowd Rejection Sampling in SFT
This subsection demonstrates the practicality of CCE by showcasing its extensive application in SFT. Rejection sampling has been proven an effective augmentation technique for SFT (Yuan et al., 2023; Zhu et al., 2023b). In a typical rejection sampling framework, given the task instruction and $k$ generated responses, low-quality responses are filtered out, and the remaining high-quality ones are then used for fine-tuning. Traditionally, the Vanilla LLM-as-a-Judge selects the best response by comparing responses in pairs and choosing the one that wins most often. In contrast, CCE naturally adapts to the scenario that rejection sampling involves more than two responses, and we refer to it as crowd rejection sampling. During pairwise comparing any two candidate responses, we effectively utilize the additional $k-2$ responses as crowd responses as introduced in Subsection 3.2. After producing crowd judgments, it ensures a more detailed and consistent judgment. We validate the crowd rejection sampling in our subsequent experiment (in Table 3), where the integration of crowd responses consistently leads to more reliable and interpretable sampling, ultimately improving the overall performance of the fine-tuned model.
4 Experiments
4.1 Experimental Setup
We conduct a comprehensive evaluation of CCE across three tasks: testing preference benchmarks, judge distillation, and SFT rejection sampling.
Model Reward Bench HelpSteer2 MTBench Human Judge Bench EvalBias Avg.
GPT-4o
Vanilla 85.2 66.1 82.1 66.3 68.5 73.6
LongPrompt 86.9 67.3 81.8 63.5 70.5 74.0
EvalPlan 88.7 65.5 81.4 62.9 74.4 74.6
16-Criteria 87.3 69.1 82.8 66.6 73.7 75.9
Maj@16 87.9 68.9 82.4 68.6 75.5 76.7
Agg@16 88.1 68.7 82.6 67.2 77.9 76.9
CCE-random@16 91.2 69.5 83.1 68.9 80.1 78.6
CCE@16 91.8 70.6 83.6 70.4 85.0 80.3
Qwen 2.5 7B-Instruct
Vanilla 78.2 60.7 76.1 58.3 57.4 66.1
CCE@16 80.4 64.2 76.7 64.0 79.4 72.9
Qwen 2.5 32B-Instruct
Vanilla 87.4 72.3 79.0 68.9 71.1 75.7
CCE@16 90.8 72.1 82.1 70.6 80.5 79.2
Qwen 2.5 72B-Instruct
Vanilla 85.2 69.5 79.5 68.3 68.5 74.0
CCE@16 93.7 68.5 88.9 75.7 85.9 82.7
Llama 3.3 70B-Instruct
Vanilla 86.4 70.4 81.1 67.1 70.6 75.1
CCE@16 91.7 71.3 83.5 69.7 79.2 79.1
Table 1: Accuracy of LLM-as-a-Judge on pair-wise comparison benchmarks. CCE can consistently enhance the LLM-as-a-Judge’s performance across 5 benchmarks, especially considerably outperforming other scaling inference strategies, like maj@16. The highest values are bolded. Here, CCE-random refers to replacing the “Criticizing Selection$+$Outcome-Removal Processing” with “Random Selection”.
Preference Benchmarks and Baselines.
We adopt 5 preference benchmarks to test LLM-as-a-Judge, including RewardBench (Lambert et al., 2024), HelpSteer2 (Wang et al., 2024d), MTBench-Human (Zheng et al., 2023), JudgeBench (Tan et al., 2025), and EvalBias (Park et al., 2024). These benchmarks provide general instructions across a wide range of tasks with diverse responses and use accuracy to measure their evaluation performance. They each focus on different aspects. For example, RewardBench covers a wider range of scenarios, while EvalBias focuses on various bias scenarios. We verify the generality of CCE on 5 LLMs and compare it against multiple baselines. In particular, we consider Vanilla, which uses the general LLM-as-a-Judge prompt implemented by RewardBench; Maj@16, where we independently judge a case 16 times and take a majority vote of the outcomes; Agg@16, where instead of majority voting, the 16 individual judgments are fed back into the LLM to aggregate a final decision; 16-Criteria, which incorporates 16 criteria with corresponding descriptions in the prompt as designed in Hu et al. (2024b) and Wang et al. (2024d); LongPrompt, where the LLM is explicitly directed to produce a longer CoT; and EvalPlan, in which an unconstrained evaluation plan is first generated based on the target case and then executed to derive the final judgment (Saha et al., 2025). Additional details on the preference benchmarks and baselines can be found in Appendix B.
Distilling CoT for Training Judge.
We start with a large preference dataset and evaluate it using the Vanilla LLM-as-a-Judge and CCE under GPT-4o-as-a-Judge, producing two CoTs. We then pair each CoT with the original preference data to form two separate training sets, which we use to fine-tune a smaller LLM as a judge. The resulting judges’ performance clearly reflects the quality and effectiveness of each CoT. We use TULU3-preference data as the distillation query while the preference benchmarks for evaluating the judge remain the same as previously introduced. Details of the training implementation are provided in Appendix C.
SFT Rejection Sampling.
Firstly, we generate a pool of 4 responses based on a given task instruction to serve as the rejection sampling base. We compare Crowd Rejection Sampling against Random Selection and a Vanilla Rejection Sampling method to select the best response for fine-tuning.
We select two datasets of different scales, LIMA (Zhou et al., 2023) ($1$K) and TULU3-SFT (Lambert et al., 2025) (sample $10$K), as instruction query. GPT-4o served as the judge LLM, while Llama-3.1-8B and Qwen-2.5-7B are used as base models for SFT. We then evaluate the generative ability of finetuned models using MTBench and AlpacaEval-2 (Dubois et al., 2024b). Details of the implementation are provided in Appendix D.
Model # of Training Samples RewardBench HelpSteer2 MTBench Human JudgeBench EvalBias Avg.
JudgeLM-7B (Zhu et al., 2023a) 100,000 46.4 60.1 64.1 32.6 42.4 49.1
PandaLM-7B (Wang et al., 2024c) 300,000 45.7 57.6 75.0 36.0 27.0 48.3
Auto-J-13B (Li et al., 2024b) 4,396 47.5 65.1 75.2 50.9 16.5 51.0
Prometheus-7B (Kim et al., 2024a) 100,000 34.6 30.8 52.8 9.3 11.7 27.8
Prometheus-2-7B (Kim et al., 2024b) 300,000 43.7 37.6 55.0 39.4 39.8 43.1
Llama-3.1-8B-Tuned
Synthetic Judgment from Vanilla 10,000 66.8 56.0 71.6 60.1 34.2 57.7
Synthetic Judgment from Vanilla 30,000 72.5 58.6 73.9 50.4 46.2 60.3
Synthetic Judgment from CCE 10,000 69.7 58.6 72.7 66.4 38.7 61.2
Synthetic Judgment from CCE 30,000 70.0 60.1 74.3 50.3 50.7 61.1
Qwen 2.5-7B-Tuned
Synthetic Judgment from Vanilla 10,000 68.1 55.6 70.7 50.2 38.4 56.6
Synthetic Judgment from Vanilla 30,000 71.4 56.2 75.1 48.2 54.7 61.1
Synthetic Judgment from CCE 10,000 68.8 56.7 71.3 49.8 40.2 57.4
Synthetic Judgment from CCE 30,000 73.3 59.5 74.9 50.1 57.1 63.0
Mix Synthetic Judgment from CCE&Vanilla 60,000 74.1 60.7 76.6 61.6 60.6 66.7
Table 2: Accuracy of Trained small LLM-as-a-Judge on pair-wise comparison benchmarks. Under the same preference pairs data, the model trained with judgments synthesized using CCE achieves more reliable evaluation results. The highest values are bolded, and the second highest is underlined.
4.2 Experiment Result
In this section, we present our main results. The preference benchmark results are shown in Table 1, the efficacy of distilling CoT for training smaller judges is summarized in Table 2, and the training efficiency of SFT rejection sampling is reported in Table 3. These three objectives are concluded across various judge LLMs and downstream tasks. Our findings for each task are as follows.
Performance on Preference Benchmarks.
Table 1 highlights CCE consistently achieves state-of-the-art performance across all preference benchmarks. First, it outperforms the Vanilla LLM-as-a-Judge, which already demonstrates reasonable reliability on multiple LLMs and benchmarks. Notably, with Qwen 2.5-72B-Instruct as the judge, our method achieves an $8.5$ increase on RewardBench and an overall average gain of $8.7$.
Second, CCE proves considerably more effective than common scaling strategies such as Maj@16 and 16-Criteria. Even with random selection, Maj@16 underperforms CCE by an average of 1.9. Although EvalPlan offers a more response-aware reasoning process than 16-Criteria, its effectiveness remains lower $2.0$-$3.7$ than CCE. Simply generating longer CoT also falls short, indicating that scaling inference-time computation calls for a more nuanced approach.
Rejection Sampling Method MTBench AlpacaEval-2
Llama 3.1 8B Base
Instructions from LIMA # 1K
Random Sampling 4.33 2.89/3.29
Vanilla Rejection Sampling 4.28 2.91/3.29
Crowd Rejection Sampling 4.53 3.02/3.31
Instructions from Tulu 3 # 10K
Random Sampling 7.51 12.81/12.45
Vanilla Rejection Sampling 7.56 19.92/17.17
Crowd Rejection Sampling 7.63 22.23/19.74
Qwen 2.5 7B Base
Instructions from LIMA # 1K
Random Sampling 8.06 14.52/9.40
Vanilla Rejection Sampling 7.91 14.40/9.44
Crowd Rejection Sampling 8.63 14.86/9.59
Instructions from Tulu 3 # 10K
Random Sampling 8.36 21.39/13.68
Vanilla Rejection Sampling 8.46 22.71/16.44
Crowd Rejection Sampling 8.41 23.78/17.56
Table 3: SFT Rejection Sampling Performance on the Instruction-Following Benchmark. The model fine-tuned with responses sampled using CCE demonstrates improved generative performance.
Strategy # of Selection Samples RewardBench HelpSteer2 MTBench Human JudgeBench EvalBias Avg.
Random-Selection 8 91.0 69.9 82.6 68.7 78.4 78.1
Praising-Selection 8 86.6 64.2 81.5 67.1 77.7 75.4
Criticizing-Selection 8 91.2 69.2 83.0 68.9 79.1 78.3
Balanced-Selection 8 90.7 68.6 82.8 67.4 78.7 77.6
Outcome-Removal Random-Selection 8 91.5 69.9 83.0 69.4 79.5 78.7
Outcome-Removal Criticizing-Selection (Sota) 8 91.5 70.1 83.2 69.5 79.9 78.8
Random-Selection 16 91.2 69.5 83.1 68.9 80.1 78.6
Praising-Selection 16 87.0 68.4 82.0 67.1 77.9 76.5
Criticizing-Selection 16 90.8 69.7 83.0 69.6 82.9 79.2
Balanced-Selection 16 90.6 69.3 82.9 68.0 79.6 78.1
Outcome-Removal Random-Selection 16 91.7 69.7 83.2 70.0 81.5 79.2
Outcome-Removal Criticizing-Selection(Sota) 16 91.8 70.6 83.6 70.4 85.0 80.3
Table 4: Accuracy of LLM-as-a-Judge on pair-wise comparison benchmarks. CCE can consistently enhance the LLM-as-a-Judge’s performance during the test-time inference phase, especially considerably outperforming maj@16 and 16-Criteria; under the same preference pairs data, the model trained with judgments synthesized using CCE achieves more reliable evaluation results.
Figure 3: Evaluation performance under scaling crowd judgments in the context. As the number of crowd judgments grows, both accuracy and CoT length generally increase.
Finally, CCE not only excels on RewardBench, the most general benchmark, but also outperforms alternatives on more challenging tasks like JudgeBench and EvalBias. Strategic crowd judgment selection further enhances performance compared to random selection. We adopt a “Criticizing Selection + Outcome Removal” strategy for our SOTA selection & processing strategy, which we discuss in detail in the following analysis.
Distilling CoT for Training Smaller Judges.
Distilling preference evaluation capabilities from powerful LLMs to train smaller LLMs is a promising direction. Table 2 demonstrates that higher-quality CoT leads to more effective distillation, resulting in improved performance for smaller judge models. Fine-tuning small models (e.g., Llama 3.1-8B and Qwen 2.5-7B) on the CoTs generated by CCE yields higher accuracy on all five benchmarks than using Vanilla CoTs. For instance, Qwen 2.5-7B trained on CCE’s synthetic CoT judgments achieves up to 73.3% on RewardBench, surpassing Vanilla baseline by a notable margin of 1.9. Moreover, combining both Vanilla and CCE synthetic judgments further boosts performance, reaching 74.1% on RewardBench and 60.6% on EvalBias. This result suggests integrating diverse CoT can further enhance accuracy and generalization.
LLM-as-a-Judge can develop biases in various scenarios, such as favoring more verbose answers. This issue is particularly pronounced in smaller judge models. As shown in Table 2, even after fine-tuning on over 100K samples, many baseline models struggle to exceed 50% accuracy. This highlights the persistent challenge of evaluation bias. Higher-quality and more comprehensive CoT distillation enhances the debiasing ability of smaller judge models. These findings suggest that many biases stem from the model focusing on limited aspects of the responses rather than assessing them holistically.
Efficacy in SFT Rejection Sampling.
As we can see in Table 3, Crowd Rejection Sampling proves effectiveness for both $1$K and $10$K data sizes, consistently yielding better finetuning performances for two base LLMs. CCE selects higher-quality responses compared to both Random Sampling and Vanilla Rejection Sampling, leading to consistent improvements in downstream instruction-following benchmarks on MTBench and AlpacaEval-2. For instance, with Llama 3.1-8B and the TULU3-SFT instructions, the fine-tuned model sees performance gains of up to $22.23$/$19.74$ on AlpacaEval-2, compared to $19.92$/$17.17$ under the Vanilla Rejection Sampling. This underscores the reliability of CCE in identifying higher-quality training examples.
Overall, the experiments confirm the flexibility and effectiveness of CCE in three key general scenarios. By leveraging crowd-based context, scaling inference-time computation, and strategically guiding the CoT process, CCE delivers consistent improvements over strong baselines.
4.3 Analysis Experiments
In this section, we conduct an in-depth analysis of the two core components of our method: crowd judgment selection & processing strategies, as well as inference scaling. We then directly examine whether the generated CoT is more comprehensive and provides a more detailed analysis of the responses under evaluation.
Selection & Processing Strategy.
We compare Random Selection, Criticizing Selection, Praising Selection, and Balanced Selection. As shown in Table 4, Criticizing Selection yields the best results, followed by Balanced Selection, while Praising Selection performs even worse than Random Selection. This suggests that lose-based judgments provide deeper insights into A/B comparisons, making criticism more informative. Additionally, the Outcome-Removal post-processing strategy substantially improves evaluation reliability, likely because final verdicts lack valuable details while introducing biases into LLM decision-making.
Inference Scaling.
Figure 3 illustrates our analysis of how scaling crowd judgments influence evaluation outcomes. Measuring accuracy and the average token length of the CoT, three preference benchmarks are tested across different judgment counts and then averaged for an overall assessment. The implementation details are in Appendix E.
As shown in Figure 3, both performance and output length generally increase as crowd judgments rise from 0 to 16. RewardBench displays a clear upward trend, while HelpSteer2 dips briefly at 2 judgments before recovering. Averaging across benchmarks (rightmost panel) confirms that more crowd judgments lead to higher accuracy and longer CoT, consistent with the inference scaling observed in studies (Brown et al., 2024; Snell et al., 2025). Furthermore, we reexamine the Table 1 and find that scaling test-time inference is a promising strategy for LLM-as-a-Judge, as demonstrated by GPT-4o-as-a-Judge. This is especially evident in bias scenarios, where the Vanilla struggles, while scaling-inference-based baselines, including CCE, show substantial gains.
Figure 4: CoT Comparison. CCE’s CoT consistently yields a higher average number of key points and a higher coverage rate across all benchmarks.
CoT Comparison.
To more directly assess whether the CoTs generated by CCE are more comprehensive than those of the Vanilla approach, we perform two analyses: Key Points Counting and Coverage Rate. First, we use GPT-4o to parse and summarize each CoT, counting the key points to measure how thoroughly the CoT is. Second, we leverage the cross-attention mechanism from Bart-base to quantify the coverage rate–how thoroughly CoT covers details in the candidate responses. We introduce the details in the Appendix F.2.
As shown in Figure 4, CCE outperforms Vanilla across all benchmarks in key point counting and coverage rate. More key points indicate that our CoT examines the text from multiple angles, while a higher coverage rate reflects a more detailed analysis. These results demonstrate that CCE offers a deeper and wider evaluation than Vanilla.
Instruction: Classify this tweet from the categories of positive, negative, or neutral. Tweet: I’m so thankful for all the wonderful people in my life
Response A
    The sentiment of this social media comment is positive. The use of the word "grateful" implies appreciation and thankfulness, while "amazing experiences" suggest that these experiences have been very fulfilling or enjoyable.
Response B
    Positive
Vanilla LLM-as-a-Judge:
    Assistant A’s . . . . . . It highlights specific words and phrases in the tweet, such as “thankful” and “wonderful people,” to justify the positive sentiment classification. This explanation adds depth and clarity to the response, making it more helpful and informative. Assistant B, . . . . . . the response less informative and helpful . . . . . . . . . - Final Verdict: [[A]]
CCE:
    AC Judgment: However, Assistant A makes a mistake by referencing words not present in the tweet, such as "grateful" and "amazing experiences," . . . . . . Assistant C also classifies the tweet as positive and provides a detailed explanation . . . . . . . . . Assistant A . . . . . , but it inaccurately references words not present in the tweet, such as "grateful" and "amazing experiences." This detracts from the accuracy of the response and could potentially confuse the user. . . . . . . Assistant B is concise and correctly classifies the tweet as positive. However, it lacks any explanation or reasoning, which limits its helpfulness and depth. . . . . . . In comparing the two, Given the importance of accuracy and explanation in sentiment analysis, . . . . . . - Final Verdict: [[B]]
Table 5: A pairwise comparison case evaluated by different methods. Preference refers to right result and Preference refers to wrong result. We emphasize the noisy evaluation elements in orange, while highlighting the useful elements of the evaluation in limongreen.
Case Study.
Table 5 presents a representative case. The vanilla is misled by fake information in Response A, causing it to overlook the Instruction and mistakenly rate Response A as more helpful. In contrast, the crowd judgment correctly identifies the error in Response A and informs subsequent evaluations. Additionally, our method produces a more detailed CoT thereby enriching the overall evaluation process, as evidenced by statements like “Assistant A does provide a brief explanation”.
5 Conclusion
In this work, we tackle the shortcomings of LLM-as-a-Judge, which stem from CoT reasoning lacking comprehensiveness and detail, by drawing inspiration from human evaluative behavior. We introduce a novel crowd-based comparative evaluation framework that enriches the CoT process to unlock more comprehensive and reliable evaluations. By scaling inference more effectively, our method serves as an efficient alternative to traditional majority voting and criteria expansion. Importantly, we demonstrate that high-quality CoT judgments boost evaluation reliability and distilling efficiency across multiple benchmarks, while broadening the scope of crowd-based evaluation applications.
Limitations
Progressive Self-Iteration Paradigm.
A limitation of our work is that we do not explore self-iteration in this study, despite its potential for enhancing the evaluation process. Our method inherently allows for iterative refinement, which could be further extended into a progressive paradigm. We leave this direction for future work, aiming to investigate how iterative self-improvement can further enhance evaluation quality and robustness.
Selection based on LLMs.
We identify that the quality of crowd judgments influences the CoT and explore a simple yet efficient selection strategy. We generate crowd responses using many LLMs, but we do not explore which LLM’s crowd response has a greater influence on crowd judgment.
Appendix A Prompt Template
We provide the prompt we used in this work for the experiment, as depicted in Figure 5. For Vanilla LLM-as-a-Judge (Figure 6), we deployed the prompt designed in MTBench, which is widely deployed in many works, e.g., RewardBench. Notably, HelpSteer2 specializes in 5 aspects, so we replace the MTBench’s aspects with these aspects when we test the method in HelpSteer2. Furthermore, we also present the prompts of baselines: LongPrompt(Figure 8) forces the CoT as long as possible; 16-Criteria (Figure 7) incorporates 16 criteria and corresponding descriptions, which are designed in Hu et al. (2024b) and Wang et al. (2024d).
Figure 5: Prompt of Our Method.
Figure 6: Prompt of Vanilla LLM-as-a-Judge.
Figure 7: Prompt of 16-Criteria LLM-as-a-Judge.
Figure 8: Prompt of LongPrompt LLM-as-a-Judge.
Appendix B Testing Preference Benchmark
B.1 Preference Benchmarks
As shown in Table 6, we give a brief introduction to preference benchmarks. Each of these benchmarks has its own strengths; thoroughly testing all of them and averaging the results is a reliable way to evaluate the method. Notably, we randomly sampled $1$K cases from the training split of EvalBias since the size of the test split is 80 items, which is too small.
Benchmarks Size Focus
RewardBench $2,985$ It covers multiple scenarios, including Chat, Chat-Hard, Safety, and Reasoning.
HelpSteeer2 $519$ It provides multiple fine-grained dimensions for evaluation, like Helpfulness, Coherence, Correctness, Complexity, Verbosity.
MTBench Human $2,665$ It provides multi-turn conversation for evaluation, and we filter the samples whose outcome is “Tie”.
JudgeBench $350$ It focuses on challenging response pairs spanning knowledge, reasoning, math, and coding
EvalBias $1,000$ It tests the robustness of judges on various scenarios containing evaluation biases.
Table 6: The brief description of Preference Benchmarks for testing.
B.2 The Implementation of Generating Crowd Judgments
To generate crowd judgments, we produce a wide range of diverse responses. We employed several API-accessible and open-source LLMs to generate these responses based on the given instructions. Since diversity is crucial, we did not limit ourselves to only the most powerful models. Specifically, we used the following LLMs: Qwen-2.5-0.5B-Instruct, Qwen-2.5-1.5B-Instruct, Qwen-2.5-3B-Instruct, Qwen-2.5-7B-Instruct, Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Mistral-Nemo10-Instruct-2407, Mistral-7B-Instruct, GPT-4o-mini, and GPT-4o. Additionally, we applied two temperature settings ($0.7$ and $1.0$) for each model. In principle, greater diversity in models and temperature configurations leads to improved performance.
Based on these crowd responses, we deployed the vanilla LLM-as-a-Judge to judge each crowd response with candidate response A/B separately using the judge LLM.
B.3 The Implementation of Baselines
For maj@16 and agg@16, we modify the temperature setting to $1.0$ to promote more diversified responses. For other inferences in baselines, we set a unified temperature as $0.1$.
B.4 The Implementation of Selection and Processing
For the selection strategy, we adopted “Criticize Selection” by choosing the crowd judgment where the outcome indicates that response A/B loses. For “Outcome Removal Processing,” we used GPT-4o-mini to eliminate the outcome segment from the judgment with a temperature of 00. The prompt is:
You are a helpful assistant. Specifically, I will provide you with the text quality judgment from an LLM-as-a-Judge evaluation of the responses from two AI assistants to an instruction. I need you to remove the final conclusion segments and only remain the evaluation analysis segments as soon as possible. ONLY OUTPUT the processed judgment.
*Judgment:* {judgment}”
B.5 The Implementation of Inference
We tested our method on multiple LLMs-as-Judge, including GPT-4o (2024-08-06), Qwen 2.5-7B-Instruct, Qwen 2.5-32B-Instruct, Qwen 2.5-72B-Instruct, and Llama 3.3-70B-Instruct. We found that reliability and consistency of evaluation can be balanced when temperature$=0.1$.
Appendix C Distilling CoT for Training Judge
C.1 Distilling Preference Source
We chose the TULU3-Preference-Mixture  [^1] https://huggingface.co/datasets/allenai/llama-3.1-tulu-3-8b-preference-mixture as the preference data source. Specifically, we prompt the LLM-as-a-Judge to generate a CoT using the given instruction along with the chosen-rejected response pairs as input. Additionally, we experiment with two training sizes: random samples of $10$K and $30$K examples.
Distilling Inference.
We use the GPT-4o as the Judge to produce the CoT, and the temperature setting is $0.1$.
C.2 The Implementation of Training Judge
Base Models.
To verify the generality of our method in Distilling CoT, we fine-tuned the preference data and corresponding CoT judgment in base LLMs: Qwen 2.5-7B-Base and Llama 3.1-8B-Base.
Training Setting.
We trained the Base LLM with a context length$=4,096$, epochs$=3$, batch size$=128$,and learning rate$=2e^{-5}$.
Appendix D SFT Data Selection
D.1 Synthetic Response Pool for Selection
To enhance the challenge and realism of the SFT Data Selection, we chose four LLMs with similar general generation capabilities as the base models for synthesizing responses. These are: GPT-4o, DeepSeek-v3, Claude-3.5-Sonnet, and Qwen 2.5-72B-Instruct. For inference, we set the temperature parameter to 0.7. We generate four responses for each instruction to serve as the basis for subsequent selection. The base instruction queries we used are two pools: LIMA and TULU3-SFT. LIMA  [^2] https://huggingface.co/datasets/GAIR/lima contains 1,000 instructions, which are regarded as high-quality; TULU3-SFT  [^3] https://huggingface.co/datasets/allenai/tulu-3-sft-mixture contains $93.9$K instruction-response pairs, and we randomly sampled $10$K instructions as the query. The latter is the latest released multilingual dataset.
D.2 The Implementation of Rejection Sampling
Under the vanilla LLM-as-a-Judge approach, we perform pairwise comparisons among four responses, awarding a score of $+1$ to the winner of each matchup. After all comparisons, the response with the highest total score is selected. Building on this, our method incorporates the remaining two responses as “crowd responses” during each evaluation, allowing us to gather additional crowd judgments.
Base Judge Model.
The base judge model is GPT-4o, and the temperature is set as $0.1$.
D.3 The Implementation of Training SFT
Base Models.
To verify the generality of our method in SFT data selection, we fine-tuned the instruction and selected response in base LLMs: Qwen 2.5-7B-Base and Llama 3.1-8B-Base.
Training Setting.
We followed the common setup for SFT, with a context length$=2048$, epochs$=3$, batch size$=128$,and learning rate$=2e^{-5}$.
Appendix E Inference Scaling
The “Vanilla” setup has no crowd judgments, “1” includes a single judgment, and even-numbered settings split judgments evenly between A and B. We use GPT-4o as the judge and sample three times per setting to obtain the average result.
Appendix F CoT Comparison
F.1 Key Points Extraction
We use the Key points statistic to measure the richness of the CoT. Firstly, we use the GPT-4o-mini to summarize the CoT to aspects and corresponding sub-points. The summarization prompt is
“Extract the key evaluation aspects and detailed points mentioned in the text below. List the aspects and points in a strictly structured format:”
“Example Input: ‘The response is accurate but lacks creativity. It includes factual details but misses key arguments.’ ”
“Example Dictionary Output:” “- Aspect: Accuracy ” “ - Sub-point: Includes factual details ” “ - Sub-point: Misses key arguments ” “- Aspect: Creativity ” “ - Sub-point: Lacks originality”
“**Input**:”
When we generate the summarized dictionary parsed output, we can get the total number of key points of each CoT.
F.2 Coverage Rate Compuataion
An attention-based approach computes mapping weights linking output tokens to input tokens. Interpretability research (Bibal et al., 2022; Vig, 2019) uses these weights to assess which input tokens influence the output. Our goal is to quantify how thoroughly CoT evaluates details in the target text, and attention-based computation provides a precise method for doing so.
Naturally, we used the bart-base  [^4] https://huggingface.co/facebook/bart-base to compute the cross-attention between the target text and the generated CoT. We extract the cross-attention weights from the last layer of the decoder. By averaging these weights across attention heads and applying a threshold$=0.3$, it calculates a coverage rate—the fraction of the target text’s tokens whose attention is above the threshold from the CoT.

### v2

@@
 Our analysis confirms that CoTs generated by ours are more comprehensive and of higher quality, and evaluation accuracy improves as inference scales.
+Our code is available at https://github.com/Don-Joey/CCE.git.
 Crowd Comparative Reasoning: Unlocking Comprehensive Evaluations for LLM-as-a-Judge
@@
 1 Introduction
+Figure 1: An overview of our method.
+By evaluating the candidate responses A/B alongside the crowd responses, the resulting crowd judgment can be used as context to enrich the evaluation of A/B responses, leading to a more comprehensive CoT judgment.
 With the prohibitive cost and limited scalability of human evaluation, LLM-as-a-Judge has emerged as a scalable framework for auto-evaluation (Chang et al., 2024; Li et al., 2024a, 2025).
@@
 Therefore, enhancing the richness and comprehensiveness of CoT reasoning is essential to improve LLM-as-a-Judge.
-Figure 1: An overview of our method.
-By evaluating the candidate responses A/B alongside the crowd responses, the resulting crowd judgment can be used as context to enrich the evaluation of A/B responses, leading to a more comprehensive CoT judgment.
 Two commonly adopted strategies aim to address this issue: majority voting (Zhang et al., 2024; Mahan et al., 2024; DeepSeek-AI, 2024) and criteria expansion (Kim et al., 2024a; Liu et al., 2024; Hu et al., 2024a).
@@
 Based on this, CCE prompts the LLM-as-a-Judge to perform a more comprehensive and deeper overall CoT judgment.
-CCE achieves a remarkable average improvement of $6.7\%$ across five evaluation benchmarks, including RewardBench, HelpSteer2, MTBench Human, JudgeBench and EvalBias.
-When applied to judge distillation, we find high-quality long CoT judgments produced by CCE have a higher efficiency for training a smaller judge model, especially enhancing bias robustness.
-Moreover, we extend CCE naturally to SFT rejection sampling, refer to crowd rejection sampling, where our approach serves as a quality signal to identify training-efficient samples from the response pool.
-This extension highlights the reliability and practicality of CCE.
-Finally, analysis experiments confirm that CCE promotes more comprehensive and deeper CoT reasoning and scales inference effectively.
+CCE achieves a remarkable average improvement of $6.7\%$ across five judge benchmarks, including RewardBench, HelpSteer2, MTBench Human, JudgeBench and EvalBias.
+When applied to judge distillation, we find that the high-quality long CoT judgments generated by CCE enable a smaller judge model to achieve higher accuracy, yielding an average improvement of $4.5\%$-$5.6\%$ (in Qwen 2.5-7B), particularly enhancing bias robustness.
+Moreover, we extend CCE naturally to SFT rejection sampling, referred to as crowd rejection sampling, where our approach serves as a quality signal to identify training-efficient samples from the response pool.
+Our enhanced rejection strategy consistently outperforms both random sampling and vanilla rejection sampling on MTBench and AlpacaEval-v2, demonstrating the reliability and practical utility of CCE in LLM alignment.
+Finally, our analysis confirms that CCE scales inference effectively and produced CoTs consistently yield more key points and capture finer-grained details within responses compared to Vanilla LLM-as-a-Judge, facilitating more comprehensive and deeper CoT reasoning.
 2 Related Work
@@
 Outcome-Removal Criticizing-Selection(Sota) 16 91.8 70.6 83.6 70.4 85.0 80.3
-Table 4: Accuracy of LLM-as-a-Judge on pair-wise comparison benchmarks.
-CCE can consistently enhance the LLM-as-a-Judge’s performance during the test-time inference phase, especially considerably outperforming maj@16 and 16-Criteria; under the same preference pairs data, the model trained with judgments synthesized using CCE achieves more reliable evaluation results.
+Table 4: Accuracy of CCE using different selection strategies on LLM-as-a-Judge benchmarks.
+Our proposed Outcome-Removal Criticizing-Selection consistently surpasses performances using other selection strategies during the test-time inference phase.
 Figure 3: Evaluation performance under scaling crowd judgments in the context. As the number of crowd judgments grows, both accuracy and CoT length generally increase.